OpenStudy (anonymous):

Solve the equation Three x minus six equals fifteen for x. A)x equals three B)x equals seven C)x equals twenty-seven D)x equals sixty-three

OpenStudy (triciaal):

3x -6 = 15 add 6 to each side 3x -6 + 6 = 15 + 6 3x = 31 divide each side by 3 3x/3 = 21/3 x = 7

OpenStudy (triciaal):

typo 3x =21

OpenStudy (anonymous):

x = 3 x = 7 x = 27 x = 63